# EmbodiedSAM: Online Segment Any 3D Thing in Real Time

### v1

@@ 3.2 Efficient Online Query Merging
Once lifting 2D masks $M_{t}^{2d}$ to accurate 3D masks $M_{t}^{cur}$, we then merge $M_{t}^{cur}$ to the previous instance masks $M_{t-1}^{pre}$ to acquire $M_{t}^{pre}$. Note when $t=1$ we have $M_{1}^{pre}=M_{1}^{cur}$ as an initialization. However, the mainstream solution for merging instance masks in previous works [33; 34; 16; 21; 17] is to traverse over all masks in $M_{t}^{cur}$ and compare each mask in $M_{t}^{cur}$ with all previous masks in $M_{t-1}^{pre}$. This process is very slow, because in order to accurately decide wheter a new mask shoule be merged into a previous mask, the geometric similarity such as mask-IoU or CD-distance is computed on the point clouds of the two masks. The computation of similarity involves all the points in each mask, which has high computation complexity. What is worse, above operations are hard to be computed in parallel, since the number of points in each mask is different and we need to pick out point clouds of each instance according to the mask one by one. To this end, we propose to represent each mask in fixed-size vectors and compute the similarity with efficient matrix operation.
@@
Second, for contrastive similarity, we aim to learn an instance-specific representation where features from the same instance should be pulled together and otherwise pushed away. This representation can be learned by contrastive training between two adjacent frames: we use MLP to map the query features $Q_{t}$ to contrastive feature $f_{t}$. Then for an instance $i$ appears in the $t$-th and $(t+1)$-th frames, we choose the two features of this instance $(f_{t}^{i},f_{t+1}^{i})$ as the position pair, and sample features from other instances $(f_{t}^{i},f_{t+1}^{k})$ as the negative pair. The detailed loss function is shown in the next subsection.
Finally, for semantic similarity, we simply adopt a MLP to predict per-category probability distribution $S\in\mathbb{R}^{K}$, where $K$ is the number of pre-defined categories. There are also other choices for this task. For example, if we adopt semantic-SAM [15] instead of SAM, we can directly utilize the semantic predictions for the 2D masks to serve as $S$ for the corresponding queries.
@@ 4 Experiment
In this section, we first describe our datasets and implementation details. Then we compare our method with state-of-the-art VFM-assisted 3D instance segmentation methods online 3D segmentation methods to validate its effectiveness. We also apply ESAM in open-vocabulary and data-efficient setting to demonstrate its application potential. Finally we conduct ablation studies to provide a comprehensive analysis on our design.
4.1 Benchmark and Implementation Details
We evaluate our method on four datasets: ScanNet [6], ScanNet200 [24], SceneNN [10] and 3RScan [30]. ScanNet contains 1513 scanned scenes, out of which we use 1201 sequences for training and the rest 312 for testing. ScanNet200 provides more fine-grained annotations on the scenes of ScanNet, which contains more than 200 categories. SceneNN contains 50 high-quality scanned scenes with instance and semantic labels. Following [32], we select 12 clean sequences for testing. 3RScan is a more challenging indoor dataset where the RGB-D sequences are acquired by fast-moving cameras. We choose its test split for testing, which contains 46 scenes. Each dataset provide both posed RGB-D sequences and reconstructed point clouds with labels.
Benchmarks: We compare different methods on four benchmarks. First, we compare with VFM-assisted 3D instance segmentation methods in Table 1. We train different methods on ScanNet200 training set (if needed) and evaluate them on ScanNet200 validation set in a class-agnostic manner. For offline methods, the input of each scene is a reconstructed point cloud and a RGB-D video, where predictions are made on the reconstructed point clouds. For online methods, the input is a streaming RGB-D video, and we map the final predicted results on $S_{t}$ to the reconstructed point clouds with nearest neighbor interpolation for comparison.
Table 1: Class-agnostic 3D instance segmentation results of different methods on ScanNet200 dataset. Following [34], we compare with conventional clustering methods and VFM-assisted 3D scene perception methods. The unit of Speed is $ms$ per frame, where the speed of VFM and other parts are reported separately.
@@ 4.1 Benchmark and Implementation Details
HDBSCAN [19] Offline 1.6 5.5 32.1
Nunes et al. [22] Offline 2.3 7.3 30.5
Felzenszwalb et al. [7] Offline 5.0 12.7 38.9
UnScene3D [25] Offline DINO [2] 15.9 32.2 58.5
SAMPro3D [31] Offline SAM [12] 18.0 32.8 56.1
SAI3D [34] Offline SemanticSAM [15] 30.8 50.5 70.6
SAM3D [33] Online SAM 20.2 35.7 55.5 1369+1518
ESAM Online SAM 37.9 58.8 75.0 1369+80
ESAM-E Online FastSAM [39] 35.9 56.3 74.0 20+80
@@
ESAM Online 26.6 46.2 63.1 10.3 23.6 50.7
ESAM-E Online 23.4 43.0 60.0 10.2 22.4 48.5
@@
TD3D [13] Offline 46.2 71.1 81.3
Oneformer3D [14] Offline 59.3 78.8 86.7
INS-Conv [16] Online 57.4
TD3D-MA [32] Online 39.0 60.5 71.3 26.0 42.8 59.2 3.5
ESAM-E Online 38.4 57.7 72.9 27.3 42.1 56.4 10.0
ESAM-E+FF [26] Online 40.8 58.9 75.7 30.2 48.6 63.6 9.8
@@
We also compare with online 3D instance segmentation methods in Table 3. Following previous works [16; 32], we train different methods on ScanNet training set and evaluate them on ScanNet validate set and SceneNN.
Finally, we evaluate the open-vocabulary 3D instance segmentation ability on the 198 categories of ScanNet200 in Table 4. Since ESAM outputs class-agnostic 3D masks, there are two methods to extend it to open-vocabulary 3D segmentation. The first is to feed the class-agnostic 3D masks to open-vocabulary mask classification model like OpenMask3D [29], which is adopted in the code of SAI3D [34]. The second is to adopt open-vocabulary 2D segmentation model to acquire the category labels for each 2D mask. Since there is one-to-one correspondence between 3D mask and 2D mask in ESAM, we can acquire the category labels for each 3D mask accordingly. Here we follow SAI3D to adopt the first method and compare with it.
@@
Implementation details: Following [32], we train ESAM in two stages. First we train a single-view perception model on ScanNet(200)-25k, which is a subset of ScanNet(200) with individual RGB-D frames, without memory-based adapters and losses for the three auxiliary tasks. Next we finetune the single-view perception model on RGB-D sequences with the adapters and full losses. To reduce memory footprint, we randomly sample 8 adjacent RGB-D frames for each scene at every iteration. In terms of hyperparameters, we set $\varphi=0.5$, $\epsilon=1.75$, $\tau=0.02$, $\alpha=0.5$ and $\beta=0.5$.
4.2 Comparison with State-of-the-art
We compare our method with the top-performance VFM-assisted 3D instance segmentation methods and online 3D instance segmentation methods as described above. We provide three versions of ESAM, namely ESAM, ESAM-E and ESAM-E+FF. ESAM adopts SAM as the VFM while ESAM-E adopts FastSAM [39] to achieve real-time inference. ESAM-E+FF not only adopts the 2D masks from FastSAM, but also fuses image features extracted by FastSAM backbone to point clouds following [26]. We also include some visualization results in for qualitative evaluation.
In terms of generalization ability, ESAM also demonstrates great performance. As shown in Table 2, when directly transferred to other datasets, ESAM still achieves leading accuracy compared with zero-shot methods. We note SAI3D even performs worse than SAM3D on 3RScan dataset, this is because it highly relies clean reconstructed 3D meshes with accurately aligned RGB frames. While in 3RScan, the camera is moving fast and thus the RGB images and camera poses are blurry.
We visualize the predictions of the above methods on ScanNet200, as shown in Figure 4. ESAM can predict accurate and fine-grained 3D instance segmentation masks, while being able to process streaming RGB-D video in real time. We also provide an online visualization to further demonstrate the practicability of ESAM in Figure 5. More details can be viewed in our video demo.
As shown in Table 3 and Table 4, ESAM also achieves state-of-the-art performance compared with previous online 3D instance segmentation methods and open-vocabulary 3D instance segmentation methods.
4.3 Analysis of ESAM
Table 5: Performance of ESAM when trained with partial training set.
Proportion AP AP50 AP25
100% 37.9 58.8 75.0
50% 37.0 58.4 75.4
20% 34.4 55.8 74.2
Data-efficient learning. We reduce the training samples by using only 20% or 50% training set and report the class-agnostic performance of ESAM on ScanNet200 in Table 5. It is shown that the performance degradation of ESAM is not significant even with limited training data. This is because 2D VFM has already provided a good initialization, thus the learning part of ESAM is easy to converge.
Decomposition of inference time. We decompose the inference time of ESAM excluding VFM in Table 6. The temporal-aware backbone consists of a sparse convolutional U-Net and several memory-based adapters. The merging process consists of similarity computation, bipartite matching and mask / representation updating. Due to the efficient design, the decoder and merging operation of ESAM only take a small proportion of inference time.
Ablation study. We first conduct ablation studies on ESAM to validate the effectiveness of the proposed methods. For architecture design, we conduct experiments on ScanNet200-25k and report class-agnostic AP and average inference latency (ms) of each frame excluding SAM in Table 4.3. It can be seen that geometric-aware pooling boosts the performance up to 1.4% while brings negligible computational overhead. Note that the prediction error on single views will accumulate on the whole scenes, so a high AP on ScanNet200-25k contributes a lot to the final performance. We can also observe that the dual-level design in ESAM achieves comparable accuracy compared with the time-consuming $F=F_{P}$ strategy, while only slightly increases the latency compared with the fully-superpoint $F=F_{S}$ strategy. For the merging strategies, we compare different design on ScanNet200 with AP reported, as shown in Table 4.3. It is shown that each auxiliary task is important for the quality of mask merging. We notice that the geometric similarity has the most significant influence on the final performance. This is because most mask pairs can be excluded based on distance.
Backbone Decoder Merging Total
3D-Unet Adapters Similarity Matching Updating
41.0 28.0 5.0 0.7 0.3 5.0 80
Table 7: Effects of the architecture design.
Method AP Latency
Replace $\mathcal{G}$ with average pooling 56.3 43.6
Set $F=F_{S}$ only 45.6 43.1
Set $F=F_{P}$ only 58.0 51.7
The final model 57.7 45.4
Table 8: Effects of the merging strategies.
Method AP
Remove box representation 28.7
Remove contrastive representation 31.6
Remove semantic representation 34.8
The final model 37.9
5 Concluding Remark
In this work, we presented ESAM, an efficient framework that leverages vision foundation models for online, real-time, fine-grained, generalized and open-vocabulary 3D instance segmentation. We propose to lift the 2D masks generated by VFM to 3D queries with geometric-aware pooling, which is followed by a dual-path query decoder to refine the queries and generate accurate 3D instance masks. Then with the query-mask correspondence, we design three auxiliary tasks to represent each 3D mask in three discriminative vectors, which enables fast mask merging with matrix operations. Extensive experimental results on four datasets demonstrates that ESAM achieves leading performance, online and real-time inference and strong generalization ability. ESAM also shows great potential in open-vocabulary and data-efficient setting. We believe ESAM brings a new paradigm on how to effectively leverage 2D VFM for embodied perception.

### v3

@@ 3.2 Efficient Online Query Merging
Once lifting 2D masks $M_{t}^{2d}$ to accurate 3D masks $M_{t}^{cur}$, we then merge $M_{t}^{cur}$ to the previous instance masks $M_{t-1}^{pre}$ to acquire $M_{t}^{pre}$. Note when $t=1$ we have $M_{1}^{pre}=M_{1}^{cur}$ as an initialization.
However, the mainstream solution for merging instance masks in previous works Yang et al. (2023); Yin et al. (2024); Liu et al. (2022); Narita et al. (2019); Lu et al. (2023) is to traverse over all masks in $M_{t}^{cur}$ and compare each mask in $M_{t}^{cur}$ with all previous masks in $M_{t-1}^{pre}$. This process is very slow, because in order to accurately decide wheter a new mask shoule be merged into a previous mask, the geometric similarity such as mask-IoU or CD-distance is computed on the point clouds of the two masks. The computation of similarity involves all the points in each mask, which has high computation complexity. What is worse, above operations are hard to be computed in parallel, since the number of points in each mask is different and we need to pick out point clouds of each instance according to the mask one by one. To this end, we propose to represent each mask in fixed-size vectors and compute the similarity with efficient matrix operation.
@@
Second, for contrastive similarity, we aim to learn an instance-specific representation where features from the same instance should be pulled together and otherwise pushed away. This representation can be learned by contrastive training between two adjacent frames: we use MLP to map the query features $Q_{t}$ to contrastive feature $f_{t}$. Then for an instance $i$ appears in the $t$-th and $(t+1)$-th frames, we choose the two features of this instance $(f_{t}^{i},f_{t+1}^{i})$ as the positive pair, and sample features from other instances $(f_{t}^{i},f_{t+1}^{k})$ as the negative pair. The detailed loss function is shown in the next subsection.
Finally, for semantic similarity, we simply adopt a MLP to predict per-category probability distribution $S\in\mathbb{R}^{K}$, where $K$ is the number of pre-defined categories. There are also other choices for this task. For example, if we adopt semantic-SAM Li et al. (2023) instead of SAM, we can directly utilize the semantic predictions for the 2D masks to serve as $S$ for the corresponding queries.
@@ 4 Experiment
Table 1: Class-agnostic 3D instance segmentation results of different methods on ScanNet200 dataset. The unit of Speed is $ms$ per frame, where the speed of VFM and other parts are reported separately.
@@
SAMPro3D Offline SAM 18.0 32.8 56.1
Open3DIS Offline GroundedSAM 34.6 43.1 48.5
SAI3D Offline SemanticSAM 28.2 47.2 67.9
SAM3D Online SAM 20.2 35.7 55.5 1369+1518
ESAM Online SAM 42.2 63.7 79.6 1369+80
ESAM-E Online FastSAM 43.4 65.4 80.9 20+80
@@
ESAM Online 28.8 52.2 69.3 14.1 31.2 59.6
ESAM-E Online 28.6 50.4 71.0 13.9 29.4 58.8
In this section, we first describe our datasets and implementation details. Then we compare our method with state-of-the-art VFM-assisted 3D instance segmentation methods online 3D segmentation methods to validate its effectiveness. We also apply ESAM in data-efficient setting to demonstrate its application potential. Finally we conduct ablation studies to provide a comprehensive analysis on our design. More supplementary experiments can be found in appendix.
4.1 Benchmark and Implementation Details
We evaluate our method on four datasets: ScanNet Dai et al. (2017), ScanNet200 Rozenberszki et al. (2022), SceneNN Hua et al. (2016) and 3RScan Wald et al. (2019). ScanNet contains 1513 scanned scenes, out of which we use 1201 sequences for training and the rest 312 for testing. ScanNet200 provides more fine-grained annotations on the scenes of ScanNet, which contains more than 200 categories. SceneNN contains 50 high-quality scanned scenes with instance and semantic labels. Following Xu et al. (2024), we select 12 clean sequences for testing. 3RScan is a more challenging indoor dataset where the RGB-D sequences are acquired by fast-moving cameras. We choose its test split for testing, which contains 46 scenes. Each dataset provide both posed RGB-D sequences and reconstructed point clouds with labels.
Benchmarks: We compare different methods on four benchmarks. First, we compare with VFM-assisted 3D instance segmentation methods in Table 1. We train different methods on ScanNet200 training set (if needed) and evaluate them on ScanNet200 validation set in a class-agnostic manner. We also train and evaluate the methods in the same way on ScanNet, where the results are put in appendix (Table 7). For offline methods, the input of each scene is a reconstructed point cloud and a RGB-D video, where predictions are made on the reconstructed point clouds. For online methods, the input is a streaming RGB-D video, and we map the final predicted results on $S_{t}$ to the reconstructed point clouds with nearest neighbor interpolation for comparison.
@@ 4.1 Benchmark and Implementation Details
TD3D Offline 46.2 71.1 81.3
Oneformer3D Offline 59.3 78.8 86.7
INS-Conv Online 57.4
TD3D-MA Online 39.0 60.5 71.3 26.0 42.8 59.2 3.5
ESAM-E Online 41.6 60.1 75.6 27.5 48.7 64.6 10.0
ESAM-E+FF Online 42.6 61.9 77.1 33.3 53.6 62.5 9.8
@@
Finally, we compare with online 3D instance segmentation methods in Table 3. Following previous works Liu et al. (2022); Xu et al. (2024), we train different methods on ScanNet training set and evaluate them on ScanNet validate set and SceneNN.
Compared methods: On Table 1 and Table 2, we compare ESAM with SAM3D Yang et al. (2023), SAMPro3D Xu et al. (2023), Open3DIS Nguyen et al. (2024) and SAI3D Yin et al. (2024). We adopt the 2D version of Open3DIS for a fair comparison. Because its 3D version makes predictions directly on reconstructed point clouds, while other methods only make predictions on RGB-D frames. On Table 3, we compare with TD3D Kolodiazhnyi et al. (2024a), Oneformer3D Kolodiazhnyi et al. (2024b), INS-Conv Liu et al. (2022) and TD3D-MA Xu et al. (2024). In terms of VFM, above methods mainly adopt SAM Kirillov et al. (2023), GroundedSAM Ren et al. (2024), SemanticSAM Li et al. (2023) and FastSAM Zhao et al. (2023).
Implementation details: Following Xu et al. (2024), we train ESAM in two stages. First we train a single-view perception model on ScanNet(200)-25k, which is a subset of ScanNet(200) with individual RGB-D frames, without memory-based adapters and losses for the three auxiliary tasks. Next we finetune the single-view perception model on RGB-D sequences with the adapters and full losses. To reduce memory footprint, we randomly sample 8 adjacent RGB-D frames for each scene at every iteration. For hyperparameters, we set $\varphi=0.5$, $\epsilon=1.75$, $\tau=0.02$, $\alpha=0.5$ and $\beta=0.5$. In the dual-level query decoder, we actually set $F=F_{S}$ for the first two iterations of mask prediction, and then set $F=F_{P}$. This smoothens the mask generation process with curriculum learning.
4.2 Comparison with State-of-the-art
We compare our method with the top-performance VFM-assisted 3D instance segmentation methods and online 3D instance segmentation methods as described above. We provide three versions of ESAM, namely ESAM, ESAM-E and ESAM-E+FF. ESAM adopts SAM as the VFM while ESAM-E adopts FastSAM to achieve real-time inference. ESAM-E+FF not only adopts the 2D masks from FastSAM, but also fuses image features extracted by FastSAM backbone to point clouds following Rukhovich et al. (2023). We also include some visualization results in for qualitative evaluation.
In terms of generalization ability, ESAM also demonstrates great performance. As shown in Table 2, when directly transferred to other datasets, ESAM still achieves state-of-the-art accuracy compared with zero-shot methods. We note offline methods perform worse on 3RScan dataset, this is because they highly rely on clean reconstructed 3D meshes with accurately aligned RGB frames. While in 3RScan, the camera is moving fast and thus the RGB images and camera poses are blurry.
We visualize the predictions of the above methods on ScanNet200, as shown in Figure 4. ESAM can predict accurate and fine-grained 3D instance segmentation masks, while being able to process streaming RGB-D video in real time. We also provide an online visualization to further demonstrate the practicability of ESAM in Figure 6 in appendix. More details can be viewed in our video demo.
As shown in Table 3, ESAM also achieves state-of-the-art performance compared with previous online 3D instance segmentation methods. Different from previous methods that only fuse 2D features to 3D point clouds, our approach utilize both 2D features and 2D masks to better guide the learning of 3D representation.
4.3 Analysis of ESAM
Table 4: Performance of ESAM when trained with partial training set.
Proportion AP AP50 AP25
100% 42.2 63.7 79.6
50% 40.2 62.3 78.4
10% 32.8 54.1 73.9
Data-efficient learning. We reduce the training samples by using only 20% or 50% training set and report the class-agnostic performance of ESAM on ScanNet200 in Table 4. It is shown that the performance degradation of ESAM is not significant with only half the training data. Moreover, ESAM still achieves state-of-the-art performance even with 10% training data (SAI3D: 28.2/47.2/67.9). This is because 2D VFM has already provided a good initialization, thus the learning part of ESAM is easy to converge.
Ablation study. We conduct ablation studies to validate the effectiveness of the proposed methods. For architecture design, we conduct experiments on ScanNet-25k and report AP and average inference latency (ms) of each frame excluding VFM in Table 4.3. It can be seen that geometric-aware pooling boosts the performance up to 1.3% while brings negligible computational overhead. Note that the prediction error on single views will accumulate on the whole scenes, so a high AP on ScanNet-25k contributes a lot to the final performance. We can also observe that the dual-level design in ESAM achieves comparable accuracy compared with the time-consuming $F=F_{P}$ strategy, while only slightly increases the latency compared with the fully-superpoint $F=F_{S}$ strategy. For the merging strategies, we compare different design on ScanNet with AP reported, as shown in Table 4.3. It is shown that each auxiliary task is important for the quality of mask merging. We notice that the geometric similarity has the most significant influence on the final performance. This is because most mask pairs can be excluded based on distance.
Table 5: Effects of the architecture design.
Method AP Latency
Replace $\mathcal{G}$ with average pooling 45.9 43.6
Set $F=F_{S}$ only 34.5 43.1
Set $F=F_{P}$ only 47.4 51.7
The final model 47.2 45.4
Table 6: Effects of the merging strategies.
Method AP
Remove box representation 33.4
Remove contrastive representation 36.9
Remove semantic representation 37.6
The final model 41.6
5 Concluding Remark
In this work, we presented ESAM, an efficient framework that leverages vision foundation models for online, real-time, fine-grained and generalized 3D instance segmentation. We propose to lift the 2D masks generated by VFM to 3D queries with geometric-aware pooling, which is followed by a dual-path query decoder to refine the queries and generate accurate 3D instance masks. Then with the query-mask correspondence, we design three auxiliary tasks to discriminatively represent each 3D mask, which enables fast mask merging with matrix operations. Extensive experimental results on four datasets demonstrates that ESAM achieves leading performance, online and real-time inference and strong generalization ability. We believe ESAM brings a new paradigm on how to effectively leverage 2D VFM for embodied perception.
Appendix A Appendix
We provide more experimental results in the appendix.
A.1 Online Visualization
We demonstrate the online 3D segmentation process in Figure 6. It is shown that ESAM can effectively merge partial segmentation results into a whole object and generate fine-grained 3D masks for the online reconstructed 3D scene.
A.2 Additional Experiments
Following Yin et al. (2024), we compare ESAM with conventional clustering methods McInnes & Healy (2017); Nunes et al. (2022); Felzenszwalb & Huttenlocher (2004); Rozenberszki et al. (2024); Caron et al. (2021) and VFM-assisted 3D scene perception methods on ScanNet in Table 7. Consistent with the results in Table 1, ESAM also achieves leading performance and speed on ScanNet in the class-agnostic 3D instance segmentation setting.
Table 7: Class-agnostic 3D instance segmentation results of different methods on ScanNet dataset. Following Yin et al. (2024), we compare with conventional clustering methods and VFM-assisted 3D scene perception methods. The unit of Speed is $ms$ per frame, where the speed of VFM and other parts are reported separately.
Method Type VFM AP AP50 AP25 Speed
HDBSCAN Offline 1.6 5.5 32.1
Nunes et al. Offline 2.3 7.3 30.5
Felzenszwalb et al. Offline 5.0 12.7 38.9
UnScene3D Offline DINO 15.9 32.2 58.5
SAMPro3D Offline SAM 16.7 31.5 57.9
Open3DIS Offline GroundedSAM 29.9 46.7 58.6
SAI3D Offline SemanticSAM 30.8 50.5 70.6
SAM3D Online SAM 20.2 34.0 53.3 1369+1518
ESAM Online SAM 48.2 70.3 85.3 1369+80
ESAM-E Online FastSAM 49.3 71.4 85.8 20+80
A.3 Analysis on Inference Time
We decompose the inference time of ESAM excluding VFM in Table 8. The temporal-aware backbone consists of a sparse convolutional U-Net and several memory-based adapters. The merging process consists of similarity computation, bipartite matching and mask / representation updating. Due to the efficient design, the decoder and merging operation of ESAM only take a small proportion of inference time. In terms of backbone, it is promising to adopt techniques like network pruning or knowledge distillation Li et al. (2016); Hou et al. (2022) to accelerate its inference speed, which we leave for future work.
Backbone Decoder Merging Total
3D-Unet Adapters Similarity Matching Updating
41.0 28.0 5.0 0.7 0.3 5.0 80
A.4 Open-vocabulary Results
@@
We further evaluate the open-vocabulary 3D instance segmentation ability on the 198 categories of ScanNet200 in Table 9. Since ESAM outputs class-agnostic 3D masks, there are two methods to extend it to open-vocabulary 3D segmentation. The first is to feed the class-agnostic 3D masks to open-vocabulary mask classification model like OpenMask3D Takmaz et al. (2023) and OpenIns3D Huang et al. (2024), which is adopted in the code of SAI3D Yin et al. (2024). The second is to adopt open-vocabulary 2D segmentation model to acquire the category labels for each 2D mask. Since there is one-to-one correspondence between 3D mask and 2D mask in ESAM, we can acquire the category labels for each 3D mask accordingly. Here we follow SAI3D to adopt the first method and compare with it.
A.5 Visualization on Geometric-aware Pooling
We provide the visualization of point-wise weights predicted in Geometric-aware Pooling in Figure 7. The points with high weights are shown in red, while ones with low weights are shown in blue. It is shown that noisy boundary with inaccurate shape information will be assigned low weight, while regions contain objects (especially small objects) are assigned high weight for better segmentation.
Figure 7: Visualization of weights in Geometric-aware Pooling. We also provide the corresponding RGB images. The points with high weights are shown in red, while ones with low weights are shown in blue.
A.6 Qualitative Results on SceneNN and 3RScan
We further visualize different methods on SceneNN and 3RScan in Figure 8. Consistent with Figure 4, ESAM predicts accurate and clean 3D masks. On the contrary, the prediction of SAM3D is noisy or incomplete while the prediction of SAI3D tends to be over-segmented.
Figure 8: Visualization results on SceneNN and 3RScan.